OpenStudy (anonymous):

In the process of completing the square, 3x2 + 7x - 12 becomes x^2+7/4x=4 true or false

sam (.sam.):

\[3x^2 + 7x - 12\] \[3(x^2+\frac{7}{3}x)-12\] \[3(x+\frac{7}{6})^2-\frac{49}{36}(3)-12\] \[3 \left(x+\frac{7}{6}\right)^2-\frac{193}{12}\]

OpenStudy (anonymous):

it would be false right @.Sam. because i got 3x^2+7x-12

sam (.sam.):

its false, another way to check this is to find the x values for both equations.

OpenStudy (anonymous):

it would be false right @.Sam. because i got 3x^2+7x-12